# Rethinking Broken Object Level Authorization Attacks Under Zero Trust Principle

### v1

@@ 6.6.1. Results
The results of evaluation on real-world BOLA vulnerabilities are shown in Table 7.
BFLA(CVE-2023-36100). The attacker obtains the userids of all users by accessing API1(GET /squareComment/getAllSquare) vertically unauthorizedly and then tampering with the userId injection point of C-API1(POST /api/User/ChangeUser Body:User,userId) to modify the information of other users. By analyzing the data flow, BolaZ found that P-API1(GET /User/GetUserInfoByid/{user_self_id}) provided the userId for C-API1. P-API1 returns the user’s own userId, so C-API1 can only use user’s own userId.
@@
Detection in the wild. We use BolaZ to scan projects in Table 3 to obtain the MSG interval of the C-API resource ID. By passing the resource ID outside the permission into the C-API parameter, we determine whether there is a vulnerability according to the response result. BolaZ reports a total of 36 vulnerabilities, of which 35 vulnerabilities have been manually confirmed as real vulnerabilities, as show in Table 8. The false positive caused by the modification of the resource ID during the propagation process. We have submitted vulnerability informations to the CVE vulnerability database. Due to ethic consideration, we will disclose the vulnerability after notifying the manufacturer to repair it. The details of BolaZ report vulnerabilities are shown in Table 9.
@@ 6.7.2. Results
The results of this comparison are summarized in Table 10. BOLARAY correctly identified 27 true BOLA vulnerabilities, all of which were also disclosed by BolaZ. However, BOLARAY missed 8 vulnerabilities reported by BolaZ because it cannot identifies BOLA vulnerabilities from SELECT statements. In contrast, BolaZ can analyzes BOLA vulnerabilities with all types of SQL statements. BOLARAY reported 3 false vulnerabilities, but they are different from those reported by BolaZ, as show in Table 11.
@@ 6.7.3. Case study of false positives
Through in-depth investigation, we have summarized the causes of these false positives as follows.
@@
Case 2: Column-level Authorization (Huang et al., 2024). Some applications require the authorization model to be more granular and limited to specific columns, leading to a false positive. In the music website project, the comment table is defined by BOLARAY as the ownership model, and a comment can only be updated/deleted by its owner. However, the comment like API (/comment/like) modifies the comment’s like_count column, which can be modified by all users.
BOLARAY reported 2 more false vulnerabilities than BolaZ. This is because BOLARAY performs authorization checking based on the artificially summarized authorization model, which will lead to failure in some application scenarios. Its authorization model does not fit the workflow of the system, and wrongly determines the context between APIs. However, BolaZ defense rules based on the system’s best-practice authorization logic solves these defects.
@@
6.8. Discussion
Threat to Validity. Through our experiments, we identified certain limitations in BolaZ’s data flow analysis when the scenarios are relatively specific. First, as systems grow more complex in functionality and logic, it becomes challenging to interpret client-side control conditions using static code analysis. BolaZ sometimes was unable to determine which resource IDs produced by P-APIs are passed to C-APIs solely by analyzing condition code, particularly in scenarios involving client-controlled deletion and modification operations.
@@ 6.8. Discussion
Future Work. First, BolaZ currently only supports databases. In the future, we plan to investigate the potential of the MSG concept in other types of storage systems. Second, improvements are needed in BolaZ’s front-end code analysis. We aim to integrate dynamic application security testing to better capture the logical relationships between P-APIs and C-APIs. Additionally, to address data flow interruptions in taint tracking, we will offer open programming interfaces, allowing users to adapt BolaZ to their system’s coding patterns through customized rules.
@@ 7.2. Static anlaysis-based BOLA Vulnerability Detection
SPACE (Near and Jackson, 2016) requires developers to provide a mapping of application resources to the basic types that occur in SPACE catalog. SPACE uses symbolic execution to extract the data exposures from source code. Then SPACE checks whether each data exposure is allowed. Cancheck (Bocic and Bultan, 2016) also requires developers to provide authorization rules for resources. RoleCast (Son et al., 2011) common software engineering patterns to determine resource authorization patterns, but RoleCast’s patterns do not apply to all web applications. MACE (Monshizadeh et al., 2014) analyzes INSERT statements to determine relationships between users and resources, but MACE only supports the ownership model (Huang et al., 2024) and cannot detect the BOLA vulnerability of SELECT statements. BOLARAY (Huang et al., 2024) combines SQL and static analysis to automatically identify BOLA vulnerabilities in database-backed applications. However, BOLARAY also requires developers to mark DAL specifications, and does not support the detection of BOLA vulnerabilities in SELECT statements.
@@ 7.3. Runtime Monitoring
There are also some works to enforce access control at system runtime. Nemesis (Dalton et al., 2009) ensures that only authenticated users can access resources under the manual authorization policy. FlowWatcher (Muthukumaran et al., 2015) found that the access control pattern of resources in most web applications is similar. Therefore, FlowWatcher requires developers to provide access policy to detect whether the policy is violated during system runtime.

### v2

@@ 6.6.1. Results
The results of the evaluation on real-world BOLA vulnerabilities are shown in Table 7.
BFLA(CVE-2023-36100). The attacker obtains the userids of all users by accessing API1(GET /squareComment/getAllSquare) vertically unauthorizedly and then tampering with the userId injection point of C-API1(POST /api/User/ChangeUser Body:User,userId) to modify the information of other users. By analyzing the data flow, BolaZ found that P-API1(GET /User/GetUserInfoByid/{user_self_id}) provided the userId for C-API1. P-API1 returns the user’s own userId, so C-API1 can only use the user’s own userId.
@@
Detection in the wild. We use BolaZ to scan projects in Table 3 to obtain the MSG interval of the C-API resource ID. By passing the resource ID outside the permission into the C-API parameter, we determine whether there is a vulnerability according to the response result. BolaZ reports a total of 36 vulnerabilities, of which 35 vulnerabilities have been manually confirmed as real vulnerabilities, as shown in Table 8. The false positive is caused by the modification of the resource ID during the propagation process. We have submitted vulnerability information to the CVE vulnerability database. Due to ethical considerations, we will disclose the vulnerability after notifying the manufacturer to repair it. The details of BolaZ report vulnerabilities are shown in Table 9.
@@ 6.7.2. Results
The results of this comparison are summarized in Table 10. BOLARAY correctly identified 27 true BOLA vulnerabilities, all of which were also disclosed by BolaZ. However, BOLARAY missed 8 vulnerabilities reported by BolaZ because it cannot identify BOLA vulnerabilities from SELECT statements. In contrast, BolaZ can analyze BOLA vulnerabilities with all types of SQL statements. BOLARAY reported 3 false vulnerabilities, but they are different from those reported by BolaZ, as shown in Table 11.
@@ 6.7.3. Case study of false positives
Through an in-depth investigation, we have summarized the causes of these false positives as follows.
@@
Case 2: Column-level Authorization (Huang et al., 2024). Some applications require the authorization model to be more granular and limited to specific columns, leading to a false positive. In the music website project, the comment table is defined by BOLARAY as the ownership model, and a comment can only be updated/deleted by its owner. However, the comment-like API (/comment/like) modifies the comment’s like_count column, which can be modified by all users.
BOLARAY reported 2 more false vulnerabilities than BolaZ. This is because BOLARAY performs authorization checking based on the artificially summarized authorization model, which will lead to failure in some application scenarios. Its authorization model does not fit the workflow of the system, and wrongly determines the context between APIs. However, BolaZ defense rules based on the system’s best-practice authorization logic solve these defects.
@@
7. Discussion and Future Work
7.1. Threat to Validity.
Through our experiments, we identified certain limitations in BolaZ’s data flow analysis when the scenarios are relatively specific. First, as systems grow more complex in functionality and logic, it becomes challenging to interpret client-side control conditions using static code analysis. BolaZ sometimes was unable to determine which resource IDs produced by P-APIs are passed to C-APIs solely by analyzing condition code, particularly in scenarios involving client-controlled deletion and modification operations.
@@ 7. Discussion and Future Work
Third, BolaZ is currently implemented on the SpringBoot framework. A fact is that the propagation logic and behavior of resource IDs remains consistent in varied languages or frameworks. Hence, BolaZ is a generic approach at the methodological level.
7.2. Future Work.
First, BolaZ currently only supports databases. In the future, we plan to investigate the potential of the MSG concept in other types of storage systems. Second, improvements are needed in BolaZ’s front-end code analysis. We aim to integrate dynamic application security testing to better capture the logical relationships between P-APIs and C-APIs. Additionally, to address data flow interruptions in taint tracking, we will offer open programming interfaces, allowing users to adapt BolaZ to their system’s coding patterns through customized rules.
@@ 8.2. Static Anlaysis-based BOLA Vulnerability Detection
SPACE (Near and Jackson, 2016) requires developers to provide a mapping of application resources to the basic types that occur in SPACE catalog. SPACE uses symbolic execution to extract the data exposures from source code. Then SPACE checks whether each data exposure is allowed. Cancheck (Bocic and Bultan, 2016) also requires developers to provide authorization rules for resources. RoleCast (Son et al., 2011) common software engineering patterns to determine resource authorization patterns, but RoleCast’s patterns do not apply to all web applications. MACE (Monshizadeh et al., 2014) analyzes INSERT statements to determine relationships between users and resources, but MACE only supports the ownership model (Huang et al., 2024) and cannot detect the BOLA vulnerability of SELECT statements. BOLARAY (Huang et al., 2024) combines SQL and static analysis to automatically identify BOLA vulnerabilities in database-backed applications. However, BOLARAY also requires developers to mark DAL specifications, and does not support the detection of BOLA vulnerabilities in SELECT statements. The principle of MOCGuard (Liu et al., 2025) and BOLARAY to detect BOLA vulnerabilities is highly similar. They are based on human-summarized authorization models and use the relationship between the database tables corresponding to the resources to infer whether the user has permission to access the corresponding resources. The four authorization models summarized by BOLARAY are more comprehensive than MOCGuard and cover more scenarios. Similarly, MOCGuard also cannot handle the dynamic authorization logic of unknown systems due to the static nature of the artificially summarized authorization model.
@@ 8.3. Runtime Monitoring
There are also some works to enforce access control at system runtime. Nemesis (Dalton et al., 2009) ensures that only authenticated users can access resources under the manual authorization policy. FlowWatcher (Muthukumaran et al., 2015) found that the access control pattern of resources in most web applications is similar. Therefore, FlowWatcher requires developers to provide an access policy to detect whether the policy is violated during system runtime.